OpenStudy (anonymous):

If 8 units represent 168 liters, what is the value of 10 units? 21 liters 160 liters 210 liters 1,344 liters

OpenStudy (knightofpentacles):

hold it

OpenStudy (knightofpentacles):

by quarts or gallons? litter amounts

OpenStudy (aaronandyson):

8 units = 168 lit 10 units = x lit 168*10/8 = x

OpenStudy (knightofpentacles):

by quarts or gallons i got a metric units :/

OpenStudy (knightofpentacles):

lol

OpenStudy (texaschic101):

make a proportion... 8/168 = 10/x --- 8 units to 168 liters = 10 units to x liters cross multiply 8x = 1680 -- divide both sides by 8 x = 1680/8 x = 210 liters

OpenStudy (anonymous):

i dont know if its quarts or gallons